find the indicated outputs for f(x) - 5x^2 -3x.
I am having difficulty in trying to figure out this problem. I thought that it is five x to the second power which means five times five which would make it 25x minus 3x but the answer comes up incorrect. Could you please help me?

Answer by rfer(16322)

It says to use the indicated outputs, what are they?
If for example x=2
f(2)=-5*2^2-3*2
f(2)=-5*4-6
f(2)=-20-6
f(2)=-26
-----
Follow that example.